OpenStudy (anonymous):

simplify the expression k^14/k^7

OpenStudy (anonymous):

there is a rule in exponents.- \[\frac{ a^b }{ a^c } => a^b-a^c = > a^(b-c)\]

OpenStudy (anonymous):

u can try to simplify the expression now ?

OpenStudy (anonymous):

lol @Gretacig no way it would be k^14-7 = k^7.

OpenStudy (anonymous):

omg :DDDDDDD my mistake :DDDD

OpenStudy (anonymous):